# Colin Rogers 1, Adriana C. Briozzo 2

### v1

@@ Abstract
A novel extension of the canonical solitonic mKdV equation is introduced which admits hybrid Ermakov-Painlevé II symmetry reduction. Application of the latter is made to obtain exact solution of Airy-type to a class of moving boundary problems of Stefan kind for this extended mKdV equation. A reciprocal transformation is then applied to the latter to generate an associated exactly solvable class of moving boundary problems for an extension of a base Casimir member of a compacton hierachy. The extended mKdV equation is shown to be embedded in a range of nonlinear evolution equations with temporal modulation as determined via the action of a class of involutory transformations with origin in Ermakov theory. Associated temporal modulation for the hybrid mKdV and KdV equation as embedded in the classical solitonic Gardner equation is delimited.
@@ 3 A Class of Moving Boundary Problems
$\frac{\partial u^{*}}{\partial t^{*}}=\frac{\partial}{\partial x^{*}}\left[\frac{\partial}{\partial x^{*}}\left(\frac{1}{u^{*}}\frac{\partial}{\partial x^{*}}\left(\frac{1}{u^{*}}\right)\right)-\frac{2}{u^{*2}}-\frac{\lambda}{3}u^{*4}(t^{*}+a)^{-2}\right],\quad x^{*}\big|_{x=0}<x^{*}<x^{*}\big|_{x=S(t)},\quad t^{*}>0,$

### v2

# Open Communications in Nonlinear Mathematical Physics ]ocnmp[ Vol.5 (2025) pp 1References Article

@@ Abstract
A novel extension of the canonical solitonic mKdV equation is introduced which admits hybrid Ermakov-Painlevé II symmetry reduction. Application of the latter is made to obtain exact solution of Airy-type to a class of moving boundary problems of Stefan kind for this extended mKdV equation. A reciprocal transformation is then applied to the latter to generate an associated exactly solvable class of moving boundary problems for an extension of a base Casimir member of a compacton hierarchy. The extended mKdV equation is shown to be embedded in a range of nonlinear evolution equations with temporal modulation as determined via the action of a class of involutory transformations with origin in Ermakov theory. Associated temporal modulation for the hybrid mKdV and KdV equation as embedded in the classical solitonic Gardner equation is delimited.
@@ 3 A Class of Moving Boundary Problems
$\tfrac{\partial u^{*}}{\partial t^{*}}=\tfrac{\partial}{\partial x^{*}}\left[\tfrac{\partial}{\partial x^{*}}\left(\tfrac{1}{u^{*}}\tfrac{\partial}{\partial x^{*}}\left(\tfrac{1}{u^{*}}\right)\right)-\tfrac{2}{u^{*2}}-\tfrac{\lambda}{3}u^{*4}(t^{*}+a)^{-2}\right],\quad x^{*}\big|_{x=0}<x^{*}<x^{*}\big|_{x=S(t)},\quad t^{*}>0,$
@@ results.
6 Conclusion
Investigation of exact solution of Stefan- type moving boundary problems for a range of canonical 1+1- dimensional solitonic equations via Painlevé II symmetry reduction as initiated in [41] has been detailed in [45], [51]-[54]. In the present work, a class of extended mKdV equations has been introduced which admit solution via hybrid Ermakov-Painlevé II symmetry reduction. Currently under investigation is exact solution of novel moving boundary problems by means of the latter procedure for certain 2+1- dimensional solitonic equations and their reciprocal associates.